OpenStudy (anonymous):

help Write the equation of a line in slope intercept form that is parallel to the line y=1/3 x+5 and passes through (-9, 5).

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

ok, what's going to be the slope of this new line?

OpenStudy (anonymous):

+y2?x1+x2 y1

OpenStudy (anonymous):

@AriPotta

OpenStudy (mathlegend):

@kendoll If it is going to parallel the slope has to be the same as the original slope.

OpenStudy (mathlegend):

So it would be still be 1/3x

OpenStudy (mathlegend):

Now, to pass through those coordinate points we need to use the point slope form. y-y1=m(x-x1) y-5=1/3(x+9) y - 5 = 1/3x + 3 +5 +5 -------------- y = 1/3x + 8

OpenStudy (anonymous):

ty :)

OpenStudy (mathlegend):

No problem